OpenStudy (anonymous):

Find the sum of the terms of the arithmetic sequence. 1, 7, 13, 19, . . ., 109 n = 19 A. 2204 B. 2090 C. 1045 D. 1102

OpenStudy (haseeb96):

solution a=1 , d=7-1=6, n=19 and the formula of A.Sequence is an=n\2{2a+(n-1)d} now put in it an=19\2{2(1)+(19-1)6} an=19\2{2+(18)6} an=19\2{2+108} an=19\2{110} since,110\2 =55 an=19(55) an=1045 this the answer i mean C is correct

OpenStudy (anonymous):

Thanks